6/18 Move PR3
- 6/19 DRMI locked.
- 6/21: Stop cooler.
- 6/20: Stop cooler
- 7/1: Open chamber
- Long term:
4/22-5/31: Yarm commissioning
- 5/27-6/30: DRMI commissioning
- 7/1-8/9: DRFPMI commissioning
- 8/12-10/4: DRFPMI ASC
- 10/7-11/15: Post Commissioning
- 11/18-12/13: Noise hunting
- Please reduce the number or speed of 16kHz channels.
Report of work
- Polarization shift was seen in PRC. Pure S pol. was confirmed at POP. Retuning beams from both ITMX and ITMY have 10% P component.
- PR gain was only 3.
- Nitrogen will be added at ETMY to increase speed of temperature up.